OpenStudy (anonymous):

need help on1st-4th derivatives of sin(x-x^2). currently stucked at 1st derivative, obtained: (sin x sin x^2)(1-2x) + (cos x cos x^2)(1-2x), but the calculator showing the 1st derivative to be: (1-2x) cos[(x-1)x]

OpenStudy (amistre64):

why you using a calculator?

OpenStudy (anonymous):

to check my 1st derivative

OpenStudy (anonymous):

its an online calculator @_@ LOL

OpenStudy (amistre64):

what D(sin) ?

OpenStudy (anonymous):

cos?

OpenStudy (amistre64):

yes, soo cos(x-x^2) * D(x-x^2) what D(x-x^2?)

OpenStudy (anonymous):

i opened sin(x-x^2) into sin x cos x^2 - cos x sin x^2 then i use product rule on both o.o or im not supposed to open it?

OpenStudy (amistre64):

ummm, i wouldnt try to invent new ways of failing just yet ...

OpenStudy (anonymous):

oooh...wait uh

OpenStudy (amistre64):

its not needed to go thru all that stuf an create alot of work for yourself

OpenStudy (anonymous):

D sin(x-x^2) = (1-2x) cos (x-x^2) ?

OpenStudy (amistre64):

D(sin(f(x))) = cos(f(x)) * D(f(x)) thats the chain rule

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

now you can distribute the cos(x-x^2) thru or leave it as is and continue to the D2

OpenStudy (anonymous):

my initial working straight continued into D2, got (1-2x)^2 [-sin(x-x^2)] -2cos(x-x^2)

OpenStudy (amistre64):

D (1-2x) cos (x-x^2) = D(1-2x) cos (x-x^2) + (1-2x) Dcos (x-x^2)

OpenStudy (amistre64):

-2cos (x-x^2) - (1-2x)^2 sin (x-x^2)

OpenStudy (amistre64):

so thats good, D3 it now

OpenStudy (anonymous):

-(1-2x)^3 cos(x-x^2) +4(1-2x) sin(x-x^2) +2(1-2x) sin(x-x^2) - cos(x-x^2) o.o

OpenStudy (amistre64):

D -2cos (x-x^2) - [ D(1-2x)^2 sin (x-x^2) + (1-2x)^2 D sin (x-x^2) ] \[2(1-2x) sin (x-x^2) - [ -4(1-2x) sin (x-x^2) + (1-2x)^3 cos (x-x^2) ]\] \[2(1-2x) sin (x-x^2) + 4(1-2x) sin (x-x^2) - (1-2x)^3 cos (x-x^2)\] \[(2-4x) sin (x-x^2) + (4-8x) sin (x-x^2) - (1-2x)^3 cos (x-x^2)\] \[(2-4x+4-8x) sin (x-x^2) - (1-2x)^3 cos (x-x^2)\] \[(6-12x) sin (x-x^2) - (1-2x)^3 cos (x-x^2)\] is what i get if i kept my head on straight :)

OpenStudy (amistre64):

I got a spurious little ^3 thats spose to be a ^2 on the end :)

OpenStudy (anonymous):

oo i check again because the main question is for the power series, so mayb its the D3 that is giving the trouble

OpenStudy (amistre64):

ugh, it right ... ^2+^1 = ^3 ...

OpenStudy (amistre64):

this just amounts to you doing alot of practice at derivitating stuff :)

OpenStudy (anonymous):

ooo D -2cos(x-x^2) is not product rule? == my head is blurrish lol

OpenStudy (anonymous):

omg ==

OpenStudy (amistre64):

it can be, but you get a 0(cos) which amounts to nothing so we pull the constant and chain the rest

OpenStudy (amistre64):

for example D 2x = D2 x + 2 Dx = 0 x + 2 1 = 0 + 2 = 2

OpenStudy (amistre64):

D 2x = 2 Dx = 2 1 = 2

OpenStudy (anonymous):

ah yes. lol. made silly mistake. currently solving d4

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

D4= 12(1-2x)^2 cos(x-x^2) + [(1-2x)^4 -12] sin(x-x^2) ==

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=4th+derivative+%28sin%28x-x%5E2%29%29 i think i see alot of the parts you got, dbl chk with that to be sure

OpenStudy (anonymous):

ooh okay. thx very much for pointing out my mistake ==

OpenStudy (anonymous):

erm the cos part, why it is (x-1)x?

OpenStudy (amistre64):

its just the way the wolf simplified it in the end ....

OpenStudy (anonymous):

owh ic... now i can get the series right... previously i dont know how on earth i can get 2lines worth of pellet for my D3... ==

OpenStudy (amistre64):

\[D: (6-12x) sin (x-x^2) - (1-2x)^3 cos (x-x^2)\] \[D(6-12x) sin (x-x^2)+(6-12x) Dsin (x-x^2) ...\]\[\hspace{10em} - [D(1-2x)^3 cos (x-x^2)+(1-2x)^3 Dcos (x-x^2)]\] \[-12sin (x-x^2)+(6-12x)(1-x^2) cos(x-x^2) ...\]\[\hspace{10em} - [-6(1-2x)^2 cos (x-x^2)-(1-2x)^4 sin (x-x^2)]\] \[-12sin (x-x^2)+(6-12x)(1-x^2) cos(x-x^2) ...\]\[\hspace{10em} + 6(1-2x)^2 cos (x-x^2)+(1-2x)^4 sin (x-x^2)\] and then whatever that simplifies to ...

OpenStudy (anonymous):

oh kay thx for your time ^.^ i can go to bed now.

OpenStudy (amistre64):

lol, sweet dreams